OpenStudy (anonymous):

suppose 60 people are asked which method is best to avoid getting the flu. the results are shown in the table below. determine the percentage for each method and complete the table. round to the nearest whole number. Method Frequency % Flu shot 15 washing hands 36 taking vitamins 7 eating healthy foods 2 Answer this Question

OpenStudy (cwrw238):

total of people is 60 so fraction saying flu shot = 15/60 as a percentage this is (15/60) * 100 = (1/4) * 100 = 25%

OpenStudy (cwrw238):

do the same processing for the other methods rounding so you get total of 100%

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw